# Multi-fidelity Bayesian Optimization: A Review

### v2

@@ 3.3 Multi-fidelity Bayesian optimization
As illustrated in Fig. 2, MF BO is the intersection of MFO and BO, and therefore it inherits the advantages of the two approaches. There are several reasons that make MF BO powerful, especially when solving engineering optimization problems:
@@
Compared with the generic BO outlined Algorithm 1, MF BO introduces two important modifications. First, it constructs in 9 a GP-based MF surrogate for the objective function using LF models and HF data. One of the most popular GP-based MF surrogates is the auto-regressive model [18]; see Section 4.2.1. Second, MF BO develops in 10 an MF acquisition function that is capable of selecting both a new design point and the fidelity level for a computational model to be called in 12. The very first example of such an acquisition function is the so-called augmented expected improvement [15]; see Section 5.2.2.
@@
  • Enhancing MF surrogates; see Section 4.
  • Innovating new MF acquisition functions based on the existing BO acquisition functions; see Section 5.
  • Applying MF BO to solving various optimization problems in science and engineering design; see Table 4.
@@ 4 Gaussian process-based multi-fidelity surrogates
Let ${\bf f}({\bf x})=[f_{1}({\bf x}),\dots,f_{T}({\bf x})]^{\intercal}$, $T\geq 2$, denote a vector of $T$ outputs of $T$ computational models with different fidelities that are used for predicting the objective function of an engineering design optimization problem. We assume that these models share the same input variables so that we do not need input variable mappings to perform information transfers between the fidelities. If the fidelities are sorted in increasing order, there exist $(T-1)$ LF predictions, and $f_{T}({\bf x})=f_{\text{H}}({\bf x})$ is the prediction of the highest-fidelity model. If there is no obvious fidelity ordering, then there is no hierarchy of the predictions.
@@
In this section, we describe popular GP-based MF surrogates constructed from $\mathcal{D}$ for use of MF BO. They include multivariate GP via linear model of coregionalization (Section 4.1), auto-regressive model and its variants (Section 4.2), graphical MF GP model (Section 4.3), Bayesian hierarchical model (Section 4.4), composition of GP models (Section 4.5), and input-augmentation GP-based MF surrogate models (Section 4.6). To obtain a structured understanding of these models, we attempt to exploit common properties shared between them.
@@ 4.1 Linear model of coregionalization
The multivariate GP via linear model of coregionalization (LMC) [see e.g., 169, 190] describes each element of ${\boldsymbol{\delta}}({\bf x})$ as a linear combination of $T$ independent zero-mean GPs with covariance functions $\kappa_{1}(\cdot,\cdot|{\boldsymbol{\phi}}_{1}),\dots,\kappa_{T}(\cdot,\cdot|{\boldsymbol{\phi}}_{T})$, where ${\boldsymbol{\phi}}_{t}$ contains hyperparameters of $\kappa_{t}$. Thus, the discrepancy vector can be written as ${\boldsymbol{\delta}}({\bf x})={\bf R}{\bf z}({\bf x})$, where ${\bf R}\in\mathbb{R}^{T\times T}$ consists of combination coefficients, and ${\bf z}(\cdot)\sim\mathcal{GP}\left({\bf 0},{\boldsymbol{\Sigma}}(\cdot,\cdot)\right)$ is a $T$-variate GP with ${\boldsymbol{\Sigma}}(\cdot,\cdot)=\operatorname{diag}\{\kappa_{1}(\cdot,\cdot|{\boldsymbol{\phi}}_{1}),\dots,\kappa_{T}(\cdot,\cdot|{\boldsymbol{\phi}}_{T})\}$. The GP prior for ${\bf f}(\cdot)$ in Eq. 4 reads
@@
Here ${\bf S}(\cdot,\cdot)={\bf R}{\boldsymbol{\Sigma}}(\cdot,\cdot){\bf R}^{\intercal}=\sum_{t=1}^{T}{\bf C}_{t}\kappa_{t}(\cdot,\cdot|{\boldsymbol{\phi}}_{t})$ is the so-called inter-group covariance matrix, where ${\bf C}_{t}={\bf r}_{t}{\bf r}_{t}^{\intercal}$ is the $t$-th coregionalization matrix and ${\bf r}_{t}$ the $t$-th column of ${\bf R}$. To determine ${\bf R}$, ${\boldsymbol{\phi}}_{t}$, and the parameters underlying ${\bf m}({\bf x})$, we condition the GP prior in Eq. 5 on the training dataset $\mathcal{D}$, which is called the training process, as detailed in Appendix B. Note that the inter-group covariance matrix ${\bf S}(\cdot,\cdot)\in\mathbb{R}^{T\times T}$ should not be confused with a significantly larger covariance matrix ${\bf K}({\bf X},{\bf X})\in\mathbb{R}^{\sum_{t=1}^{T}N_{t}\times\sum_{t=1}^{T}N_{t}}$ that is used for the training process and prediction equations, i.e., Eqs. (B.5) and (B.6) of Appendix B.
@@ 4.2.1 KOH auto-regressive model
If the fidelities are ordered, the surrogate for predicting the output of fidelity $t$ ($t=2,\dots,T$) can be constructed from the surrogate that approximates the output of fidelity $(t-1)$ using the auto-regressive model (i.e., KOH model). The KOH model reads [18]
@@
where $b_{t-1}$ in Eq. 6b is an unknown correlation coefficient and $\delta_{t}(\cdot)\sim\mathcal{GP}\left(m_{\delta,t}(\cdot),\kappa_{\delta,t}(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,t})\right)$ is the discrepancy function independent of $f_{t-1}(\cdot),\dots,f_{1}(\cdot)$. Here $b_{t-1}$ and $\delta_{t}(\cdot)$ play similar roles as $b$ and $\delta(\cdot)$ in Eq. 2, respectively. A common choice for the mean function of $\delta_{t}(\cdot)$ is $m_{\delta,t}({\bf x})=\sum_{i=1}^{p}\beta_{i}\zeta_{i}({\bf x})$, where $\beta_{i}$ and $\zeta_{i}({\bf x})$ are combination coefficients and basis functions, respectively. The covariance function $\kappa_{\delta,t}(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,t})$ is often the squared exponential covariance function [18, 10]. Other covariance functions for use of GP-based MF modeling include Matern [133] and composite covariance functions [191]. $\text{cov}\left[f_{t}({\bf x}),f_{t-1}({\bf x}^{\prime})|f_{t-1}({\bf x})\right]$ in Eq. 6c represents the covariance of two random variables $f_{t}({\bf x})$ and $f_{t-1}({\bf x}^{\prime})$ given that $f_{t-1}({\bf x})$ is known. The Markov property in Eq. 6c implies that observing $f_{t-1}({\bf x}^{\prime})$ provides no information for predicting $f_{t}({\bf x})$ if $f_{t-1}({\bf x})$ is observed.
@@
Let $z_{1}({\cdot})\sim\mathcal{GP}\left(0,\kappa_{\delta,1}(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,1})\right)$ and $z_{2}({\cdot})\sim\mathcal{GP}\left(0,\kappa_{\delta,2}(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,2})\right)$, where the covariance functions $\kappa_{\delta,1}$ and $\kappa_{\delta,2}$ are parameterized by the hyperparameter vectors ${\boldsymbol{\phi}}_{\delta,1}$ and ${\boldsymbol{\phi}}_{\delta,2}$, respectively. Eq. 7 can be rewritten as
@@
Here ${\bf I}_{2}$ denotes the $2$-by-$2$ identity matrix, ${\bf e}_{1}=[1,0]^{\intercal}$, and ${\bf e}_{2}=[0,1]^{\intercal}$.
Thus, the GP prior for ${\bf f}({\bf x})$ in Eq. 9 can be written using the form in Eq. 5 with the inter-group covariance matrix
@@
In the general case when there exist $T$ model fidelities ($T\geq 2$), the terms of Eq. 9 are
@@ 4.2.2 Hierarchical Kriging model
Han and Görtz [19] proposed a hierarchical Kriging model by making three modifications to the KOH model in Eq. 6. First, they modeled the surrogate that approximates the lowest-fidelity output $f_{1}({\bf x})$ in Eq. 6a as a sum of an unknown constant and a stationary random process. Second, they replaced $f_{t-1}({\bf x})$ in Eq. 6b with the best linear unbiased predictor of Kriging, denoted by $\mu_{\text{f},t-1}({\bf x})$ [32]. Finally, they used $z_{t}({\cdot})\sim\mathcal{GP}\left(0,\kappa_{\text{z},t}(\cdot,\cdot|{\boldsymbol{\phi}}_{\text{z},t})\right)$ instead of $\delta_{t}(\cdot)\sim\mathcal{GP}\left(m_{\delta,t}(\cdot),\kappa_{\delta,t}(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,t})\right)$, $t=1,\dots,T$. The hierarchical Kriging model reads
@@
where the unknown constant $a$ in Eq. 13a is given in Eq. 2. The GP prior for the hierarchical Kriging model can be derived from the model in Eq. 9 for ${\bf R}={\bf I}_{T}$ and $\boldsymbol{\mu}({\bf x})=[a,b_{1}\mu_{\text{f},1}({\bf x}),\cdots,b_{T-1}\mu_{\text{f},T-1}({\bf x})]^{\intercal}$.
@@ 4.2.3 Recursive model
Gratiet and Garnier [20] introduced the recursive model by making two modifications to the KOH model. First, they adopted the hybrid additive/multiplicative approach by replacing the regression coefficient $b_{t-1}$ in Eq. 6b with an adjustment coefficient function $\rho_{t-1}({\bf x})$, which is similar to $\rho({\bf x})$ in Eq. 2. Second, rather than using the GP prior $f_{t-1}({\bf x})$ in Eq. 6b, they used the GP posterior $\hat{f}_{t-1}({\bf x})$ constructed from the data $\mathcal{D}_{t-1}$ of fidelity $(t-1)$. Accordingly, the recursive model reads
@@
Here $\rho_{t-1}({\bf x})={\boldsymbol{\zeta}}_{t-1}^{\intercal}({\bf x}){\boldsymbol{\beta}}_{t-1}$, as shown in Fig. 4, is the linear combination of a finite number of basis functions, and
@@ 4.3 Graphical multi-fidelity Gaussian process
In an attempt to generalize the KOH model, Ji et al. [101] developed the graphical multi-fidelity Gaussian process (GMGP) for cases when the hierarchy of LF models remains unclear. The GMGP approach involves constructing a directed acyclic graph (DAG) where each node represents a computational model, and each directed edge connecting two nodes represents the hierarchy of their fidelities. If the graph is in a topological ordering such that every directed edge $(t^{\prime},t)$ from $t^{\prime}$ to $t$, then $t^{\prime}<t$ and the prediction of node $t^{\prime}$ is made before making the prediction of node $t$. Figure 6 shows an example of a topological-ordering DAG of four nodes.
@@
Consider node $t$ and let $t^{\prime}\in\text{Pa}(t)$ be a parent node of $t$. Thus, $t\in\text{Ch}(t^{\prime})$ is a child node of $t^{\prime}$. Let $V_{\text{s}}$ denote a set of all source nodes that have no lower-fidelity node, and $\bar{V}_{\text{s}}$ denote a set of non-source nodes. The DAG in Figure 6, for example, has $V_{\text{s}}=\{1,2\}$ and $\bar{V}_{\text{s}}=\{3,4\}$.
@@
Here ${\bf f}({\bf x})\in\mathbb{R}^{T}$ and ${\boldsymbol{\mu}}({\bf x})\in\mathbb{R}^{T}$ are described in Eqs. (12a) and (12d), respectively. ${\bf z}(\cdot)\sim\mathcal{GP}\left({\bf 0},{\boldsymbol{\Sigma}}(\cdot,\cdot)\right)$ is a $T$-variate GP with ${\boldsymbol{\Sigma}}(\cdot,\cdot)=\operatorname{diag}\{\kappa_{\delta,1}(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,1}),\dots,\kappa_{\delta,T}(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,T})\}$. The condition in Eq. 16b holds for $t^{\prime}\in\text{Pa}(t)$ or for $t,t^{\prime}\in V_{\text{s}}$, with $t\neq t^{\prime}$. The coefficient matrix ${\bf R}$ in Eq. 16a is defined as
@@
where the correlation coefficient $b_{j,t}$ relates nodes $j$ and $t$ as indicated in Fig. 6. We see that, while the GMGP model is a type of the LMC model described in Section 4.1, it is more general than the KOH model in Eq. 12 as the coefficient matrix ${\bf R}$ in Eq. 17 becomes that in Eq. 12c if the fidelities of all nodes are ordered.
@@
While the computational cost for training the GMGP in Eq. 16 is $\mathcal{O}(\sum_{t=1}^{T}N_{t})^{3}$, where $N_{t}$ is the sample size of data associated with node $t$, that for training the model in Eq. 18 is $\mathcal{O}\left(T\times\max\{N_{t}^{3},t=1,\dots,T\}\right)$. Inference in GMGP and its applications to several numerical experiments and to emulation of heavy-ion collisions are detailed in Ji et al. [101].

### v3

@@ 3.3 Multi-fidelity Bayesian optimization
As illustrated in Fig.˜2, MF BO is the intersection of MFO and BO, and therefore it inherits the advantages of the two approaches. There are several reasons that make MF BO powerful, especially when solving engineering optimization problems:
@@
Compared with the generic BO outlined in Algorithm˜1, MF BO introduces two important modifications. First, it constructs in Line 9 a GP-based MF surrogate for the objective function using LF models and HF data. One of the most popular GP-based MF surrogates is the auto-regressive model [18]; see Section˜4.2.1. Second, MF BO develops in Line 10 an MF acquisition function that is capable of selecting both a new design point and the fidelity level for a computational model to be called in Line 12. The very first example of such an acquisition function is the so-called augmented expected improvement [15]; see Section˜5.2.2.
@@
  • Enhancing MF surrogates; see Section˜4.
  • Innovating new MF acquisition functions based on the existing BO acquisition functions; see Section˜5.
  • Applying MF BO to solving various optimization problems in science and engineering design; see Table˜4.
@@ 4 Gaussian process-based multi-fidelity surrogates
Let ${\bf f}({\bf x})=[f_{1}({\bf x}),\dots,f_{T}({\bf x})]^{\intercal}$, $T\geq 2$, denote a vector of $T$ outputs of $T$ computational models with different fidelities that are used for predicting the objective function of an engineering design optimization problem. We assume that these models share the same input variables so that we do not need input variable mappings to perform information transfers between the fidelities. If the fidelities are sorted in increasing order, there exist (T1)(T-1) LF predictions, and $f_{T}({\bf x})=f_{\text{H}}({\bf x})$ is the prediction of the highest-fidelity model. If there is no obvious fidelity ordering, then there is no hierarchy of the predictions.
@@
In this section, we describe popular GP-based MF surrogates constructed from $\mathcal{D}$ for use of MF BO. They include multivariate GP via linear model of coregionalization (Section˜4.1), auto-regressive model and its variants (Section˜4.2), graphical MF GP model (Section˜4.3), Bayesian hierarchical model (Section˜4.4), composition of GP models (Section˜4.5), and input-augmentation GP-based MF surrogate models (Section˜4.6). To obtain a structured understanding of these models, we attempt to exploit common properties shared between them.
@@ 4.1 Linear model of coregionalization
The multivariate GP via linear model of coregionalization (LMC) [see e.g., 169, 190] describes each element of ${\boldsymbol{\delta}}({\bf x})$ as a linear combination of $T$ independent zero-mean GPs with covariance functions $\kappa_{1}(\cdot,\cdot|{\boldsymbol{\phi}}_{1}),\dots,\kappa_{T}(\cdot,\cdot|{\boldsymbol{\phi}}_{T})$, where ${\boldsymbol{\phi}}_{t}$ contains hyperparameters of $\kappa_{t}$. Thus, the discrepancy vector can be written as ${\boldsymbol{\delta}}({\bf x})={\bf R}{\bf z}({\bf x})$, where ${\bf R}\in\mathbb{R}^{T\times T}$ consists of combination coefficients, and ${\bf z}(\cdot)\sim\mathcal{GP}\left({\bf 0},{\boldsymbol{\Sigma}}(\cdot,\cdot)\right)$ is a $T$-variate GP with ${\boldsymbol{\Sigma}}(\cdot,\cdot)=\operatorname{diag}\{\kappa_{1}(\cdot,\cdot|{\boldsymbol{\phi}}_{1}),\dots,\kappa_{T}(\cdot,\cdot|{\boldsymbol{\phi}}_{T})\}$. The GP prior for ${\bf f}(\cdot)$ in Eq.˜4 reads
@@
Here ${\bf S}(\cdot,\cdot)={\bf R}{\boldsymbol{\Sigma}}(\cdot,\cdot){\bf R}^{\intercal}=\sum_{t=1}^{T}{\bf C}_{t}\kappa_{t}(\cdot,\cdot|{\boldsymbol{\phi}}_{t})$ is the so-called inter-group covariance matrix, where ${\bf C}_{t}={\bf r}_{t}{\bf r}_{t}^{\intercal}$ is the $t$-th coregionalization matrix and ${\bf r}_{t}$ the $t$-th column of $\bf R$}. To determine $\bf R$}, ${\boldsymbol{\phi}}_{t}$, and the parameters underlying ${\bf m}({\bf x})$, we condition the GP prior in Eq.˜5 on the training dataset $\mathcal{D}$, which is called the training process, as detailed in Appendix B. Note that the inter-group covariance matrix ${\bf S}(\cdot,\cdot)\in\mathbb{R}^{T\times T}$ should not be confused with a significantly larger covariance matrix ${\bf K}({\bf X},{\bf X})\in\mathbb{R}^{\sum_{t=1}^{T}N_{t}\times\sum_{t=1}^{T}N_{t}}$ that is used for the training process and prediction equations, i.e., Eqs. (B.5) and (B.6) of Appendix B.
@@ 4.2.1 KOH auto-regressive model
If the fidelities are ordered, the surrogate for predicting the output of fidelity $t$ ($t=2,\dots,T$) can be constructed from the surrogate that approximates the output of fidelity (t1)(t-1) using the auto-regressive model (i.e., KOH model). The KOH model reads [18]
@@
where $b_{t-1}$ in Eq.˜6b is an unknown correlation coefficient and $\delta_{t}(\cdot)\sim\mathcal{GP}\left(m_{\delta,t}(\cdot),\kappa_{\delta,t}(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,t})\right)$ is the discrepancy function independent of $f_{t-1}(\cdot),\dots,f_{1}(\cdot)$. Here $b_{t-1}$ and $\delta_{t}(\cdot)$ play similar roles as $b$ and $\delta(\cdot)$ in Eq.˜2, respectively. A common choice for the mean function of $\delta_{t}(\cdot)$ is $m_{\delta,t}({\bf x})=\sum_{i=1}^{p}\beta_{i}\zeta_{i}({\bf x})$, where $\beta_{i}$ and $\zeta_{i}({\bf x})$ are combination coefficients and basis functions, respectively. The covariance function $\kappa_{\delta,t}(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,t})$ is often the squared exponential covariance function [18, 10]. Other covariance functions for use of GP-based MF modeling include Matern [133] and composite covariance functions [191]. $\text{cov}\left[f_{t}({\bf x}),f_{t-1}({\bf x}^{\prime})|f_{t-1}({\bf x})\right]$ in Eq.˜6c represents the covariance of two random variables $f_{t}({\bf x})$ and $f_{t-1}({\bf x}^{\prime})$ given that $f_{t-1}({\bf x})$ is known. The Markov property in Eq.˜6c implies that observing $f_{t-1}({\bf x}^{\prime})$ provides no information for predicting $f_{t}({\bf x})$ if $f_{t-1}({\bf x})$ is observed.
@@
Let $z_{1}({\cdot})\sim\mathcal{GP}\left(0,\kappa_{\delta,1}(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,1})\right)$ and $z_{2}({\cdot})\sim\mathcal{GP}\left(0,\kappa_{\delta,2}(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,2})\right)$, where the covariance functions $\kappa_{\delta,1}$ and $\kappa_{\delta,2}$ are parameterized by the hyperparameter vectors ${\boldsymbol{\phi}}_{\delta,1}$ and ${\boldsymbol{\phi}}_{\delta,2}$, respectively. Eq.˜7 can be rewritten as
@@
Here ${\bf I}_{2}$ denotes the 22-by-22 identity matrix, ${\bf e}_{1}=[1,0]^{\intercal}$, and ${\bf e}_{2}=[0,1]^{\intercal}$.
Thus, the GP prior for ${\bf f}({\bf x})$ in Eq.˜9 can be written using the form in Eq.˜5 with the inter-group covariance matrix
@@
In the general case when there exist $T$ model fidelities ($T\geq 2$), the terms of Eq.˜9 are
@@ 4.2.2 Hierarchical Kriging model
Han and Görtz [19] proposed a hierarchical Kriging model by making three modifications to the KOH model in Eq.˜6. First, they modeled the surrogate that approximates the lowest-fidelity output $f_{1}({\bf x})$ in Eq.˜6a as a sum of an unknown constant and a stationary random process. Second, they replaced $f_{t-1}({\bf x})$ in Eq.˜6b with the best linear unbiased predictor of Kriging, denoted by $\mu_{\text{f},t-1}({\bf x})$ [32]. Finally, they used $z_{t}({\cdot})\sim\mathcal{GP}\left(0,\kappa_{\text{z},t}(\cdot,\cdot|{\boldsymbol{\phi}}_{\text{z},t})\right)$ instead of $\delta_{t}(\cdot)\sim\mathcal{GP}\left(m_{\delta,t}(\cdot),\kappa_{\delta,t}(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,t})\right)$, $t=1,\dots,T$. The hierarchical Kriging model reads
@@
where the unknown constant $a$ in Eq.˜13a is given in Eq.˜2. The GP prior for the hierarchical Kriging model can be derived from the model in Eq.˜9 for ${\bf R}={\bf I}_{T}$ and $\boldsymbol{\mu}({\bf x})=[a,b_{1}\mu_{\text{f},1}({\bf x}),\cdots,b_{T-1}\mu_{\text{f},T-1}({\bf x})]^{\intercal}$.
@@ 4.2.3 Recursive model
Gratiet and Garnier [20] introduced the recursive model by making two modifications to the KOH model. First, they adopted the hybrid additive/multiplicative approach by replacing the regression coefficient $b_{t-1}$ in Eq.˜6b with an adjustment coefficient function $\rho_{t-1}({\bf x})$, which is similar to $\rho({\bf x})$ in Eq.˜2. Second, rather than using the GP prior $f_{t-1}({\bf x})$ in Eq.˜6b, they used the GP posterior $\hat{f}_{t-1}({\bf x})$ constructed from the data $\mathcal{D}_{t-1}$ of fidelity (t1)(t-1). Accordingly, the recursive model reads
@@
Here $\rho_{t-1}({\bf x})={\boldsymbol{\zeta}}_{t-1}^{\intercal}({\bf x}){\boldsymbol{\beta}}_{t-1}$, as shown in Fig.˜4, is the linear combination of a finite number of basis functions, and
@@ 4.3 Graphical multi-fidelity Gaussian process
In an attempt to generalize the KOH model, Ji et al. [101] developed the graphical multi-fidelity Gaussian process (GMGP) for cases when the hierarchy of LF models remains unclear. The GMGP approach involves constructing a directed acyclic graph (DAG) where each node represents a computational model, and each directed edge connecting two nodes represents the hierarchy of their fidelities. If the graph is in a topological ordering such that every directed edge $(t^{\prime},t)$ from $t^{\prime}$ to $t$, then $t^{\prime}<t$ and the prediction of node $t^{\prime}$ is made before making the prediction of node $t$. Figure˜6 shows an example of a topological-ordering DAG of four nodes.
@@
Consider node $t$ and let $t^{\prime}\in\text{Pa}(t)$ be a parent node of $t$. Thus, $t\in\text{Ch}(t^{\prime})$ is a child node of $t^{\prime}$. Let $V_{\text{s}}$ denote a set of all source nodes that have no lower-fidelity node, and $\bar{V}_{\text{s}}$ denote a set of non-source nodes. The DAG in Figure˜6, for example, has $V_{\text{s}}=\{1,2\}$ and $\bar{V}_{\text{s}}=\{3,4\}$.
@@
Here ${\bf f}({\bf x})\in\mathbb{R}^{T}$ and ${\boldsymbol{\mu}}({\bf x})\in\mathbb{R}^{T}$ are described in Eqs. (12a) and (12d), respectively. ${\bf z}(\cdot)\sim\mathcal{GP}\left({\bf 0},{\boldsymbol{\Sigma}}(\cdot,\cdot)\right)$ is a $T$-variate GP with ${\boldsymbol{\Sigma}}(\cdot,\cdot)=\operatorname{diag}\{\kappa_{\delta,1}(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,1}),\dots,\kappa_{\delta,T}(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,T})\}$. The condition in Eq.˜16b holds for $t^{\prime}\in\text{Pa}(t)$ or for $t,t^{\prime}\in V_{\text{s}}$, with $t\neq t^{\prime}$. The coefficient matrix $\bf R$} in Eq.˜16a is defined as
@@
where the correlation coefficient $b_{j,t}$ relates nodes $j$ and $t$ as indicated in Fig.˜6. We see that, while the GMGP model is a type of the LMC model described in Section˜4.1, it is more general than the KOH model in Eq.˜12 as the coefficient matrix $\bf R$} in Eq.˜17 becomes that in Eq.˜12c if the fidelities of all nodes are ordered.
@@
While the computational cost for training the GMGP in Eq.˜16 is $\mathcal{O}(\sum_{t=1}^{T}N_{t})^{3}$, where $N_{t}$ is the sample size of data associated with node $t$, that for training the model in Eq.˜18 is $\mathcal{O}\left(T\times\max\{N_{t}^{3},t=1,\dots,T\}\right)$. Inference in GMGP and its applications to several numerical experiments and to emulation of heavy-ion collisions are detailed in Ji et al. [101].